# Diffusion Model in Latent Space for Medical Image Segmentation Task

### v1

@@ II Related work
Medical Image Segmentation is the task of assigning a label to each pixel to determine its class membership. This problem has been extensively studied using various AI architectures. Early models following a one-to-one approach include fully convolutional networks [1] and encoder–decoder architectures with skip connections, such as UNet [3] and UNet++ [4], which extend skip connections for improved feature propagation. Traditional CNN-based models, including nnUNet [24] and ResUNet [27], have further demonstrated strong performance in medical image segmentation. In addition, transformer-based architectures, such as SegFormer [21], have been applied for more comprehensive image analysis, and hybrid approaches combining hypernetworks, such as HyperSeg [2], have also been explored.
@@ III Method
The proposed model, MedSegLatDiff, integrates a conditional diffusion-based model with a VAE. This section first provides a brief overview of conditional diffusion models for medical image segmentation and a VAE model for encoding and decoding both medical images and segmentation masks. It then focuses on latent diffusion models (LDMs) with conditioning mechanisms operating in latent space for this task.
@@ III-A Segmentation in image pace
where $X^{(i)}\in\mathbb{R}^{H\times W\times C}$ denotes the $i$-th medical image, $S^{(i)}\in\mathbb{R}^{H\times W}$ is its corresponding segmentation mask, and $N$ is the total number of image–mask pairs in the dataset. The main idea when applying DM for segmentation task is to gradually add noise $\epsilon\sim\mathcal{N}(0,I)$ to a mask $S$ through a forward process and progressively remove noise from the noisy image via a reverse process. The forward process starts with the original image $S_{0}$ sampled from the mask distribution $p_{\textrm{mask}}(S)$ and is defined as:
@@
where $\epsilon^{\prime}\sim\mathcal{N}(0,I)$ introduces stochasticity into the reverse process, and $t=T,\dots,1$ denotes the reverse diffusion steps. Therefore, after $T$ reverse steps, the model can generate multiple segmentation masks $S_{0}$, which approximate samples drawn from the original mask distribution.
However, the setup described above only generates new masks and cannot directly perform segmentation on an input image. Specifically, given a medical image $X$ and its corresponding segmentation mask $M$, an additional channel is appended to the input to incorporate anatomical information. Therefore, the mask and the input image are concatenated channel-wise to obtain
@@ V Conclusion
In this paper, we proposed MedSegLatDiff, a conditional diffusion model operating in the latent space via VQ-VAE for medical image segmentation. We designed two VQ-VAE modules for images and masks to encode data into a compact latent representation. For perceptual mask compression, incorporating weighted cross-entropy as the reconstruction loss significantly improved the encoding–decoding quality. Our experiments show that using five samples yields stable and reliable results, enabling MedSegLatDiff to outperform previous one-to-one segmentation methods and compete effectively with recent one-to-many approaches. This strategy provides an efficient solution that improves both training and sampling efficiency, thereby enhancing segmentation performance through a sampling process that mimics a group of experts. For future work, we plan to explore more advanced conditioning mechanisms, such as classifier guidance [5], classifier-free guidance [23], and flow–based generative modeling, to further improve stability, calibration, and uncertainty estimation of the model’s outputs.

### v2

@@ II Related work
Medical Image Segmentation is the task of assigning a label to each pixel to determine its class membership. This problem has been extensively studied using various AI architectures. Early models following a one-to-one approach include fully convolutional networks [1] and encoder–decoder architectures with skip connections, such as UNet [2] and UNet++ [3], which extend skip connections for improved feature propagation. Traditional CNN-based models, including nnUNet [23] and ResUNet [31], have further demonstrated strong performance in medical image segmentation. In addition, transformer-based architectures, such as SegFormer [20], TransUNet [29], SwinUNet [30], have been applied for more comprehensive image analysis, and hybrid approaches combining hypernetworks, such as HyperSeg [28], have also been explored.
@@ III Method
The proposed model, MedSegLatDiff, integrates a conditional diffusion-based model with VAEs. This section first provides a brief overview of conditional diffusion models for medical image segmentation and VAEs model for encoding and decoding both medical images and segmentation masks. It then focuses on latent diffusion models (LDMs) with conditioning mechanisms operating in latent space for this task.
@@ III-A Segmentation in image pace
where $X^{(i)}\in\mathbb{R}^{H\times W\times C}$ denotes the $i$-th medical image, $S^{(i)}\in\{0,1\}^{H\times W}$ is its corresponding segmentation mask, and $N$ is the total number of image–mask pairs in the dataset. The main idea when applying DM for segmentation task is to gradually add noise $\epsilon\sim\mathcal{N}(0,I)$ to a mask $S$ through a forward process and progressively remove noise from the noisy image via a reverse process. The forward process starts with the original image $S_{0}$ sampled from the mask distribution $p_{\textrm{mask}}(S)$ and is defined as:
@@
where $\epsilon^{\prime}\sim\mathcal{N}(0,I)$ introduces stochasticity into the reverse process, and $t=T,\dots,1$ denotes the reverse diffusion steps. Therefore, by repeating the $T$ reverse steps, the model can generate multiple segmentation masks $S_{0}$, which approximate samples drawn from the original mask distribution.
However, the setup described above only generates arbitrary masks and cannot perform segmentation on an input image. Specifically, given a medical image XXand its corresponding segmentation mask $S$, we adapt the model for the segmentation task by appending $X$ as an additional input channel to incorporate anatomical information. Therefore, the mask and the input image are concatenated channel-wise to obtain
@@ V Conclusion
In this paper, we proposed MedSegLatDiff, a conditional diffusion model operating in the latent space via VQ-VAE for medical image segmentation. We designed two VQ-VAE modules for images and masks to encode data into a compact latent representation. For perceptual mask compression, incorporating weighted cross-entropy as the reconstruction loss significantly improved the encoding–decoding quality. Our experiments show that using five samples yields stable and reliable results, enabling MedSegLatDiff to outperform previous one-to-one segmentation methods and compete effectively with recent one-to-many approaches. This strategy provides an efficient solution that improves both training and sampling efficiency, thereby enhancing segmentation performance through a sampling process that mimics a group of experts. For future work, we plan to explore more advanced conditioning mechanisms, such as classifier guidance [4], classifier–free guidance [22], and flow–based generative modeling, to further improve stability, calibration, and uncertainty estimation of the model’s outputs.